OpenStudy (anonymous):

find the derivative of 3^(2x^2-3x)

OpenStudy (jhannybean):

\[\large d(3^{2x^2-3x})\]This follows the format: \(d(a^x) = a^x \cdot d(x)\cdot \ln(a)\)

OpenStudy (jhannybean):

So let's first find the derivative of the power. Can you solve : \(d(2x^2-3x)\)?

OpenStudy (jhannybean):

@dolphins121 ?

OpenStudy (anonymous):

uhmm

OpenStudy (anonymous):

the derivative of that would be 4x-3 right?

OpenStudy (jhannybean):

Yes, then just use the form to find your answer :)

OpenStudy (jhannybean):

So what would \(a^x\) be? and \(\ln(a)\)?

OpenStudy (anonymous):

a^x would be 3^4x-3 ?

OpenStudy (jhannybean):

That would be \(a^{d(x)}\)

OpenStudy (anonymous):

so a^x is the original problem?

OpenStudy (jhannybean):

yep :)

OpenStudy (anonymous):

and ln(a) is 3?

OpenStudy (jhannybean):

nope. \(\ln(a) = \ln(3)\)

OpenStudy (jhannybean):

because \(a=3\)

OpenStudy (anonymous):

yeah my bad thats what i meant. so the answer is 3^(4x-3) x ln(3) ?

OpenStudy (jhannybean):

nope, fit the format. \(\color{blue}{\text{Originally Posted by}}\) @Jhannybean \[\large d(3^{2x^2-3x})\]This follows the format: \(d(a^x) = a^x \cdot d(x)\cdot \ln(a)\) \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

ohh so it's 3^(2x^2-3x) ⋅ 3^(4x-3) ⋅ ln(3) ?

OpenStudy (jhannybean):

Let me refresh, it's showing question marks :(

OpenStudy (anonymous):

oh should i rewrite it?

OpenStudy (jhannybean):

Oh! you got it :) Good job. I got that too :D

OpenStudy (anonymous):

thanks to you ! :)

OpenStudy (jhannybean):

sorry about the notation as well, \(d(3^{2x^2-3x}) = \frac{d}{dx}(3^{2x^2-3x}) = (4x-3)3^{2x^2-3x}\cdot \ln(3)\)

OpenStudy (anonymous):

its np. thank you so much !

OpenStudy (jhannybean):

Good job :)

OpenStudy (anonymous):

thanks :)